OpenStudy (anonymous):

i need to solve systems by substitution, can someone walk me through this? y - 2x = -5 y - x = -3

jimthompson5910 (jim_thompson5910):

pick one equation and solve for any variable. I'm going to solve y - x = -3 for y y - x = -3 y - x+x = -3+x y = x - 3

jimthompson5910 (jim_thompson5910):

then focus on the other equation and plug y = x-3 into it y - 2x = -5 x-3 - 2x = -5 ... replace every copy of y with x-3 -x-3 = -5 ... ... ... x = ???

OpenStudy (anonymous):

so im still confused, idk where i substitute the Y

jimthompson5910 (jim_thompson5910):

what do you mean? which part?

OpenStudy (anonymous):

like i know if you solve for x then you substitute it in like 2(x-3) blah blah blah like that right?

jimthompson5910 (jim_thompson5910):

do you see how I got y = x-3 ?

OpenStudy (anonymous):

you removed it from the left side and put it on the right?

OpenStudy (anonymous):

make y=x-3 (x-3) -2x=5

jimthompson5910 (jim_thompson5910):

I isolated y in y - x = -3 to get y = x-3

OpenStudy (anonymous):

x-3-2x=5 -x-3=5 x=-8 so y=8-3 y=5

OpenStudy (anonymous):

then you substitute the y in for the next equation to get x?

OpenStudy (anonymous):

@Kamil69 ?

jimthompson5910 (jim_thompson5910):

Kamil69 it should be y = x-3 y = -8-3 y = -11

OpenStudy (anonymous):

y dosent equal 5 its -11?

OpenStudy (anonymous):

yes you are right y=-11

OpenStudy (anonymous):

now we substitute? -11 -2x = -5?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

You're done. You found x = -8 and y = -11.

OpenStudy (anonymous):

i dont know how we found x..

OpenStudy (anonymous):

because y=x-3 so y= -8-3 y=-11 so you can check your answer by y-2x=5 -11-2(-8)=5 -11+16=5 5=5

jimthompson5910 (jim_thompson5910):

Kamil69 shows you how in the steps above

OpenStudy (anonymous):

can we do a couple more together so i can figure ouut how to do this on my own?

OpenStudy (anonymous):

i have assignment do tonight, but i can help you tomorrow if you want

OpenStudy (anonymous):

2x - y = -8 2x + y = 4

OpenStudy (anonymous):

make y=4-2x

OpenStudy (anonymous):

ok i got that part

OpenStudy (anonymous):

so put back in first equation

OpenStudy (anonymous):

2x(4-2x) = -8?

OpenStudy (anonymous):

you should have 2x-(4-2x)=-8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now 2x-4+2x=-8

OpenStudy (anonymous):

-8x - 4x = -8

OpenStudy (anonymous):

no

OpenStudy (anonymous):

4x-4=-8

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

4x = -4

OpenStudy (anonymous):

x = -1?

OpenStudy (anonymous):

4x=-8+4

OpenStudy (anonymous):

4x=-4 so x=-1

OpenStudy (anonymous):

then sub it in for the next equation 2(-1) - y = -8

OpenStudy (anonymous):

no you dont have because wes have y=4-2x y=4-2(-1) y=4+2 y=6

OpenStudy (anonymous):

ok im starting to get it...one more pls? x + y =2 3x -2y =6

OpenStudy (anonymous):

i got y = -5

OpenStudy (anonymous):

-3, -5?

OpenStudy (anonymous):

@Kamil69

OpenStudy (anonymous):

no y=0 x=2